OpenStudy (anonymous):

Helpful steps please, find the phase shift of y=2cos(6x+pi).

OpenStudy (anonymous):

easy way is to set argument equal to 0 and solve for x. in your case: \[6x+\pi=0 \Rightarrow 6x=-\pi \Rightarrow x=-\frac{ \pi }{ 6 }\]

OpenStudy (anonymous):

Thanks! That helped me solve my other questions as well :)